40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1330     Entry time: Fri Feb 20 19:31:16 2009
Author: Yoichi 
Type: Update 
Category: LSC 
Subject: MICH low gain problem 
Last night, we found that MICH UGF was too low. Even after re-aligning the PDs, it was still too low.
Today, I compared the UGFs of MICH and PRC when in the DRMI configuration locked with the single demod. signals.
In this configuration, MICH signal comes from REFL33Q and the PRC signal comes from REFL33I (the same PD).
The PRC UGF was about 100Hz whereas MICH was only ~10Hz.
Since they uses the same PD, the low gain is not caused by the PD.
I checked conlog history and confirmed there is no change in the MICH->BS path in the last few days.
I also checked the svn history of chans directory for changes in filters. Nothing problematic found.

Then I noticed that the susvme computers were overloaded.
This time, I rebooted all the FE computers just in case.

Then the MICH gain was somewhat recovered (by a factor of 3 or so). Don't know why.

I adjusted the DD_handoff script to set the MICH gain to 0.7 before the bounce-roll filter is engaged.
ELOG V3.1.3-